OpenStudy (anonymous):

using multi-step equations how do u do 4/5= -4/7g

OpenStudy (gw2011):

(-4/7)g=4/5 g=(4/5)(-7/4) You can cancel the 4s g=-7/5

OpenStudy (anonymous):

we cant have fractions

OpenStudy (anonymous):

g = -1.4

OpenStudy (gw2011):

g=-7/5=-1.4

OpenStudy (anonymous):

k but i still dont get how to do it without fractions somehow have to make them into whole numbers

OpenStudy (gw2011):

The only other thing you can do is: (-4/7)g=4/5 Multiply both sides by 35 to eliminate the denominator -20g=28 g=-28/20=-7/5=-1.4

OpenStudy (anonymous):

thanks

OpenStudy (gw2011):

You're welcome